AsTEX (French) CervanTEX (Spanish) CSTUG (Czech/Slovak) CTEX (Chinese) CyrTUG (Russian) DANTE (German) DK-TUG (Danish) Estonian User Group εϕτ (Greek) GuIT (Italian) GUST (Polish) GUTenberg (French) GUTpt (Portuguese) ÍsTEX (Icelandic) ITALIC (Irish) KTUG (Korean) Lietuvos TEX’o Vartotojų Grupė (Lithuanian) MaTEX (Hungarian) Nordic TEX Group (Scandinavian) NTG (Dutch) TEXCeH (Slovenian) TEX México Tirant lo TEX (Catalan) TUG (international) TUGIndia TUG-Philippines UK TUG ViêtTUG (Vietnamese) http://www.tug.org/texcollection  2013 proTEXt: an easy to install TEX system for MS Windows: based on MiKTEX, with the TEXstudio editor front-end. TEX Live: a rich TEX system to be installed on hard disk or a portable device such as a USB stick. Comes with support for most modern systems, including GNU/Linux, Mac OS X, and Windows. MacTEX: an easy to install TEX system for Mac OS X: the full TEX Live distribution, with the TeXShop front-end and other Mac tools. CTAN: a snapshot of the Comprehensive TEX Archive Network, a set of servers worldwide making TEX software publically available. proTEXt ist ein einfach zu installierendes TEX-System für MS Windows, basierend auf MiKTEX und TEXstudio als Editor. TEX Live ist ein umfangreiches TEX-System, zur Installation auf Festplatte oder einem portablem Medium, z. B. USB-Stick. Binaries für viele Platformen sind enthalten. MacTEX ist ein einfach zu installierendes TEXSystem für Mac OS X, mit einem vollständigen TEX Live, sowie TEXShop als Frontend und weitere Programme. CTAN ist ein weltweites Netzwerk von ftp-Servern für TEX-Software. Auf der DVD befindet sich ein kompletter Abzug des deutschen CTAN-Knotens dante.ctan.org. proTEXt : un système TEX pour Windows facile à installer, basé sur MikTEX avec l’éditeur TEXstudio. TEX Live : un système TEX complet qui peut être installé sur disque dur ou en mode portable sur une clé USB. Fonctionne sur la plupart des systèmes modernes, dont GNU/Linux, Mac OS X et Windows. MacTEX : un système TEX facile à installer pour Mac OS X. Il comporte une distribution TEX Live complète ainsi que l’éditeur TeXShop et d’autres outils pour Mac. CTAN : une copie du Comprehensive TEX Archive Network, le réseau de serveurs assurant la distribution publique de TEX et ses amis dans le monde entier.  ^{^{Collection 2013} proTEXt^{TEX for MS Windows based on MiKTEX} MacTEX^{TEX for Mac OS X including full TEX Live}  TEX Live^{TEX for GNU/Linux, Unix, and MS Windows} CTAN^{Comprehensive TEX Archive Network}^{Editors: Thomas Feuerstack (proTEXt) • Karl Berry (TEX Live) Richard Koch (MacTEX) • Manfred Lotz (CTAN)} DVD_{June 2013}^{TEX}^{����� e.V._{www.dante.de} gutenberg.eu.org www.tug.org}}